Q1. The experience of a house agent indicates that he can provide suitable accomodation for
75% of the clients who come to him. If on a particular occasion, 6 clients approach him
indepndently, calculate the probability that
i. Less than 4 clients
ii. Exactly 4 clients
iii. At least 5 clients, will get satisfactory accommodation
1
Expert's answer
2021-08-29T17:33:17-0400


Let the random variable X be the number of clients approaching the house agent and get satisfactory accommodation. Then we can say that X follows a binomial distribution with p=0.75 and n=6

Given the the probability of success, p, and the number of trials, n, for a binomial distribution;

P(X=x) = "\begin{pmatrix}
 n \\
 X 
\end{pmatrix}"px(1-p)n-x for x = 0, 1,2,3,4,5 and 6


1.The probability that less than 4 clients will get satisfactory accommodation


Given the binomial distribution with p=0.75 and n=6;

P(X<4) = P(x=0) +P(x=1) + P(x=2) +P(x=3)

P(x=0) = "\begin{pmatrix}
 6 \\
 0 
\end{pmatrix}"0.750(1-0.75)6-0 = 0.0002 

P(x=1) = "\begin{pmatrix}
 6 \\
 1 
\end{pmatrix}"0.751(1-0.75)6-1 = 0.0044

P(x=2) =  "\begin{pmatrix}
 6 \\
 2 
\end{pmatrix}"0.752(1-0.75)6-2 = 0.0330

P(x=3) = "\begin{pmatrix}
 6 \\
 3 
\end{pmatrix}"0.753(1-0.75)6-3 = 0.1318

Therefore,

P(X<4) = 0.0002 + 0.0044 +0.0330 + 0.1318 = 0.1694



2.Exactly 4 clients


P(4) =  "\begin{pmatrix}
 6 \\
 4 
\end{pmatrix}"0.754(1-0.75)6-4 = 0.2966

The probability that exactly four clients will get satisfactory accommodation will be approximately equal to 0.2966


3.At least 5 clients, will get satisfactory accommodation


P(X≥5) = P(5) + P(6)

P(5) = "\begin{pmatrix}
 6 \\
 5 
\end{pmatrix}"0.755(1-0.75)6-5 = 0.3560

P(6) =  "\begin{pmatrix}
 6 \\
 6 
\end{pmatrix}"0.756(1-0.75)6-6 = 0.1780

Therefore,

P(X≥5) = P(5) + P(6) = 0.3560 + 0.1780 ≃ 0.5340